OpenStudy (anonymous):

Please help! Reduce (sin65degrees)/(1+cos65degrees) to a single function of one angle. a. tan32degrees 30' b. tan130degrees c. cot32degrees 30' d. cot130degrees

OpenStudy (ash2326):

We have \[\frac{\sin 65}{1+\cos 65}\] @IloveCharlie what's cos 0?

OpenStudy (ash2326):

@IloveCharlie are you here?

OpenStudy (anonymous):

Oh yea, just a second

OpenStudy (ash2326):

np

OpenStudy (anonymous):

1?

OpenStudy (ash2326):

Good so, I'll rewrite it as \[\frac{\sin 65}{\cos 0+\cos 65}\] Do you know what's \(\large \cos A+\cos B\)?

OpenStudy (ash2326):

@IloveCharlie ???

OpenStudy (anonymous):

No, I don't :(

OpenStudy (ash2326):

NP \[\large \cos A+\cos B=2\cos (\frac A2-\frac B2)\cos (\frac A 2+\frac B2)\] so we get now \[\large \frac{\sin 65}{2\cos (\frac 02-\frac {65}2)\cos (\frac 0 2+\frac {65}2)}\] Do you get this?

OpenStudy (anonymous):

Oh, yea, makes sense

OpenStudy (ash2326):

good...would you tell me what's \(\sin 2\theta\) in terms of \(\sin \theta\ and\ \cos \theta\)?

OpenStudy (ash2326):

@IloveCharlie ??

OpenStudy (anonymous):

Uhhhh, ? No :(

OpenStudy (ash2326):

it's given as \[\sin 2\theta=2 \sin \theta \times \cos \theta\] using this formula. we'll get \[\large \frac{2\sin (\frac{65}{2})\times\cos(\frac{65}{2})}{2\cos (-\frac {65}2)\cos (\frac {65}2)}\] what would we get now?

OpenStudy (ash2326):

@IloveCharlie ??

OpenStudy (anonymous):

K sorry, this keyboard isn't working. Just a second

OpenStudy (ash2326):

cool:)

OpenStudy (anonymous):

.9063/1.422 ? I'm thinking I did that wrong :/

OpenStudy (ash2326):

don't compute the values, just simplify

OpenStudy (anonymous):

sin(65)/1+cos(65) ?

OpenStudy (ash2326):

nope, check again. I think you have written the question again:)

OpenStudy (anonymous):

:O I'm not sure :/

OpenStudy (ash2326):

\[\large\large \frac{\cancel 2\sin (\frac{65}{2})\times\cancel{\cos(\frac{65}{2})}}{\cancel 2\cos (-\frac {65}2)\cancel{\cos (\frac {65}2)}}\] what would you get now?

OpenStudy (anonymous):

sin(65/2)/cos(-65/2)

OpenStudy (ash2326):

@IloveCharlie what's \[\large \cos (-x)\] ???

OpenStudy (anonymous):

cos(x)?

OpenStudy (ash2326):

good so we'll get now \[\large \frac{\sin \frac{65}{2}}{\cos \frac{65}{2}}\] now what would you get?

OpenStudy (anonymous):

tan65/2?

OpenStudy (ash2326):

yeah so it's \[\tan 32.5 degrees\] convert the 0.5 degrees into minutes

OpenStudy (anonymous):

So a.! THANK YOU SO MUCH ASH!

OpenStudy (ash2326):

you're welcome, did you understand?

OpenStudy (anonymous):

Yes, thanks for explaining each step

OpenStudy (ash2326):

np:D